# EfficientQAT: Efficient Quantization-Aware Training for Large Language Models

### v2

@@ 1 Introduction
Recent advancements in large language models (LLMs) (Touvron et al., 2023; Bubeck et al., 2023; Chiang et al., 2023; Xu et al., 2023a; Ying et al., 2024) have demonstrated impressive capabilities in diverse language tasks such as reasoning (Clark et al., 2018; 2019; Zellers et al., 2019), cognitive processing (Fu et al., 2023; Xu et al., 2023a), and agent-based applications (Qin et al., 2023a; b). However, these models are characterized by their extensive parameters, which pose significant challenges for memory footprint and bandwidth (Kim et al., 2023b; Xu et al., 2024a).
@@
(a) 2-bit quantization comparisons
(b) Q-PEFT comparisons
(c) training speed comparisons
Figure 1: (a) EfficientQAT significantly surpasses existing uniform quantization methods, and is either superior to or comparable with vector quantization techniques. (b) EfficientQAT markedly outperforms existing Q-PEFT methods. (c) EfficientQAT can complete the QAT of 70B models on a single NVIDIA A100-80GB GPU with less time.
To address these challenges, we introduce a novel quantization-aware training framework called EfficientQAT. This framework combines the advantages of fully trainable parameters and end-to-end training, similar to native QAT (Ma et al., 2024), while maintaining the training efficiency of PTQ (Cheng et al., 2023; Shao et al., 2023) and Q-PEFT (Xu et al., 2023b). EfficientQAT introduces block-wise training of all parameters (Block-AP) to enhance the optimizable space and mitigate quantization accuracy loss. Block-AP sequentially trains all parameters, including original full-precision weights and quantization parameters (step sizes and zero points), within each transformer block. Several works have been developed based on block-wise reconstruction. However, previous approaches focus on designing additional trainable parameters, such as clipping thresholds for OmniQuant (Shao et al., 2023), weight rounding for AutoRound (Cheng et al., 2023) and BRECQ (Li et al., 2021), or LoRA (Hu et al., 2021) parameters for CBQ (Ding et al., 2023). Our Block-AP is the first to directly train all parameters during block-wise reconstruction, achieving superior performance compared to previous methods (see Table 7). Block-AP successfully demonstrates that complex trainable parameter design is unnecessary for effective block-wise reconstruction in LLMs quantization. Furthermore, we introduce end-to-end training of quantization parameters (E2E-QP) to account for inter-block interactions. E2E-QP keeps the quantized weights fixed and trains only the quantization parameters (step sizes) end-to-end.
Thanks to the integration of the proposed Block-AP and E2E-QP, EfficientQAT characterizes itself as a fast-converging, memory-efficient, and high-performing quantization technique. For instance, EfficientQAT can obtain a 2-bit Llama-2-70B model on a single A100-80GB GPU in just 41 hours, with less than 3 points accuracy degradation on 5 zero-shot common-sense tasks compared to its full-precision counterpart (69.48 vs. 72.41). We also evaluate EfficientQAT across scenarios involving model compression and instruction-tuning. In model compression, as illustrated in Figure 1(a), EfficientQAT significantly outperforms existing uniform quantization methods by approximately 5 points on accuracy in the challenging 2-bit quantization setting. It also matches the performance of vector quantization methods (Egiazarian et al., 2024; Tseng et al., 2024) in this scenario. In terms of instruction tuning, as shown in Figure 1(b), EfficientQAT consistently outperforms existing Q-PEFT methods, including QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and PEQA (Kim et al., 2023a). For instance, EfficientQAT surpasses PEQA (Kim et al., 2023a) with 4.5 points MMLU accuracy when fine-tuning with Alpaca dataset.
@@ 2 Related Works
Post-Training Quantization of LLMs. PTQ is a pivotal technique for accelerating and deploying LLMs. Quantization approaches generally fall into two categories: weight-only quantization (Frantar et al., 2022; Dettmers et al., 2023b; Lee et al., 2023a; Kim et al., 2023b) and weight-activation quantization (Xiao et al., 2023; Liu et al., 2023c; Wei et al., 2022; 2023; Yuan et al., 2023; Zhao et al., 2023; Ashkboos et al., 2023; Li et al., 2023a; Ashkboos et al., 2024). Weight-only quantization focuses on compressing weights into low-bit formats, reducing memory demands and enhancing the efficiency of memory-bounded computations in LLMs (Lin et al., 2024; Yuan et al., 2024). Conversely, weight-activation quantization compresses both weights and activations, thus further decreasing the overhead associated with matrix multiplications (Lin et al., 2024). Recent advancements in weight-only quantization include the introduction of vector quantization methods by QUIP#Tseng et al. (2024) and AQLMEgiazarian et al. (2024). These methods have shown promising performance but also introduce significant overhead (Gong et al., 2024). Our research continues to explore uniform quantization, which is preferred for its compatibility with hardware implementations.
@@ 3.1 Method Overview
In this section, we introduce EfficientQAT, a novel quantization-aware training framework for LLMs that enhances memory efficiency. As illustrated in Figure 2, traditional QAT approaches train the full-precision weights $\mathbf{W}$ and quantization parameters $s$ (step sizes) and $z$ (zero points) simultaneously in an end-to-end manner, which significantly increases the memory requirements due to the large number of parameters involved. To address this issue, EfficientQAT adopts a two-stage strategy: block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). In the Block-AP phase, model parameters and quantization parameters are trained block-by-block using reconstruction loss, which not only allows for precise calibration with full training but also reduces memory consumption (Li et al., 2021; Shao et al., 2023) by block-wise training. Following this, the E2E-QP phase fixes the quantized weights and trains the step sizes exclusively on target datasets, thus achieving inter-block interaction in a memory-efficient way. Details on Block-AP and E2E-QP are further described in Sections 3.2 and 3.3, respectively.
@@ 3.3 End-to-End Training of Quantization Parameters
Table 1: Llama 2 & 3 average zero-shot accuracy on 5 common-sense reasoning tasks ($\uparrow$). Detailed number of each task can be found at Table 15 (4-bit), Table 16 (3-bit), and Table 17 (2-bit). “Group” indicates group size for uniform quantization and codebook scheme for vector quantization. “bold” indicates best results for uniform quantization.
Method Bits Type Group (code) 2-7 2-13 2-70 3-8 3-70
FP16 16 - 16 64.86 67.81 72.41 68.58 75.33
RTN 4 uniform 128 64.52 67.50 72.26 67.79 73.98
GPTQ 4 uniform 128 64.24 67.27 72.39 67.80 74.74
AWQ 4 uniform 128 64.54 67.61 72.44 68.24 74.77
OmniQ 4 uniform 128 64.52 67.10 72.39 - -
AutoRound 4 uniform 128 64.39 67.36 72.47 - -
QuIP# 4 vector - 64.48 67.28 72.17 - -
EfficientQAT 4 uniform 128 64.27 67.52 72.62 68.43 74.57
RTN 3 uniform 128 62.06 65.77 70.83 58.72 65.29
GPTQ 3 uniform 128 62.48 66.18 71.47 60.58 71.28
AWQ 3 uniform 128 62.82 66.14 71.41 64.82 73.65
OmniQ 3 uniform 128 62.42 66.18 71.07 - -
AutoRound 3 uniform 128 63.72 66.68 71.24 - -
QuIP# 3 vector - 63.52 66.26 72.13 - -
EfficientQAT 3 uniform 128 64.02 67.28 71.76 67.35 72.42
OmniQ 2 uniform 128 46.98 53.56 54.87 - -
AutoRound 2 uniform 128 54.50 60.72 67.70 - -
EfficientQAT 2 uniform 128 59.50 63.88 68.93 59.37 67.57
AQLM 2 vector 2x8 57.61 62.22 69.85 - -
AQLM 2 vector 1x16 61.85 64.95 70.84 64.10 70.10
QuIP# 2 vector - 60.61 64.44 70.91 - -
EfficientQAT 2 uniform 64 60.14 64.48 69.48 60.76 67.89
@@ 4.1 EfficientQAT for LLMs Quantization
Table 2: Comparison with QAT methods on Llama-2 & 3 models.
Model Method Bits Group PPL Average
Wiki C4 Accuracy
2-7B LLM-QAT 3 128 6.02 - -
2-7B BitDistiller 3 128 5.97 - -
2-7B EfficientQAT 3 128 5.81 7.34 64.02
2-7B PB-LLN 2 64 20.37 44.88 37.59
2-7B DB-LLN 2 64 7.23 9.62 55.12
2-7B EfficientQAT 2 64 6.86 8.50 60.14
2-7B LLM-QAT 2 128 9.30 - -
2-7B BitDistiller 2 128 8.08 - -
2-7B EfficientQAT 2 128 7.19 8.79 59.50
3-8B PB-LLM 2 128 24.70 79.20 38.80
3-8B DB-LLM 2 128 13.60 19.20 51.80
3-8B EfficientQAT 2 128 9.80 13.22 59.37
3-70B PB-LLM 2 128 11.60 34.50 46.00
3-70B EfficientQAT 2 128 6.07 9.23 67.57
Evaluation. We assess the zero-shot accuracy of five common-sense reasoning tasks using the v0.4.2 lm-evaluation-harness***https://github.com/EleutherAI/lm-evaluation-harness. The tasks include WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), Arc-Easy (Clark et al., 2018), and Arc-Challenge (Clark et al., 2018). We also measure the perplexity of Wikitext2 and C4 with a 2048 context length, as done in previous studies (Frantar et al., 2022; Shao et al., 2023).
@@
PTQ Baseline. We compare our results with PTQ methods from uniform quantization such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), ApiQ (Liao & Monz, 2024) and AutoRound (Cheng et al., 2023), and vector quantization including QuIP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024). Note that if a result is the best of uniform quantization, we set it to bold.
PTQ Results. The accuracy and perplexity results are presented in Table 1 and Table 3, respectively. It seems that minimal performance degradation with 4-bit group quantization, and even simple round-to-nearest (RTN) quantization achieves respectable results. Therefore, EfficientQAT offers a slight perplexity improvement ($\sim$0.02) in 4-bit uniform quantization and comparable zero-shot accuracy with existing methods. As the number of quantization bits decreases, EfficientQAT’s performance advantage increases. In 3-bit quantization, EfficientQAT surpasses the optimization-based method AutoRound by approximately 0.5% in zero-shot accuracy and outperforms OmniQuant by 0.14 to 0.43 in perplexity. Notably, in 2-bit quantization, EfficientQAT outperforms AutoRound (Cheng et al., 2023) by 5% in zero-shot accuracy for Llama-2-7b. EfficientQAT also exceeds previous vector quantization methods like AQLM (Egiazarian et al., 2024) using a 2x8 codebook. While AQLM (Egiazarian et al., 2024) with a 1x16 codebook and QuIP# (Tseng et al., 2024) perform slightly better in 2-bit scenarios, they introduce notable computational overhead (Gong et al., 2024) and can even reduce inference speed. Thus, EfficientQAT successfully pushes the limits of uniform quantization in LLMs, making hardware-friendly uniform quantization competitive with more complex vector quantization methods. However, we can also find that Llama-3 models suffer more performance degeneration after quantization (Huang et al., 2024), which may caused by the full training with 15T tokens other than 2T tokens.
QAT Baseline. We also compare our results with existing QAT methods, including LLM-QAT (Liu et al., 2023e), BitDistiller (Du et al., 2024), PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024).
QAT Results. As demonstrated in Table 2, EfficientQAT markedly surpasses other QAT methods in performance. Specifically, EfficientQAT reduces perplexity by 0.89 compared to BitDistiller (Du et al., 2024) when applied to the Llama-2-7B model with w2g128 quantization. In the more challenging scenario of Llama-3 quantization (Huang et al., 2024), EfficientQAT exceeds the performance of DB-LLM (Chen et al., 2024) by 5.98 in C4 perplexity and by 8.57 points in average accuracy with the w2g128 quantization.
Table 3: Llama 2 & 3 Wikitext2 and C4 perplexity ($\downarrow$), context length 2048.
Wikitext 2 C4
Method Bits Type Group 2-7 2-13 2-70 3-8 3-70 2-7 2-13 2-70 3-8 3-70
FP16 16 - 16 5.47 4.88 3.32 6.14 2.85 6.97 6.47 5.52 8.88 6.73
GPTQ 4 uniform 128 5.61 4.98 3.42 6.53 3.38 7.13 6.56 5.58 9.35 7.02
AWQ 4 uniform 128 5.62 4.97 3.41 6.55 3.26 7.13 6.56 5.58 9.41 6.98
OmniQ 4 uniform 128 5.58 4.95 3.40 - - 7.12 6.56 5.58 - -
QuIP# 4 vector - 5.56 4.95 3.38 - - 7.07 6.54 5.56 - -
EfficientQAT 4 uniform 128 5.53 4.93 3.39 6.47 3.17 7.07 6.54 5.58 9.26 6.94
GPTQ 3 uniform 128 6.29 5.42 3.85 9.58 5.25 7.89 7.00 5.85 11.66 8.64
AWQ 3 uniform 128 6.24 5.32 3.74 8.16 4.69 7.84 6.94 5.81 11.49 7.91
OmniQ 3 uniform 128 6.03 5.28 3.78 - - 7.75 6.98 5.85 - -
QuIP# 3 vector - 5.79 5.10 3.56 - - 7.32 6.72 5.67 - -
EfficientQAT 3 uniform 128 5.81 5.12 3.61 7.09 4.19 7.34 6.73 5.71 10.06 7.43
OmniQ 2 uniform 128 11.06 8.26 6.55 - - 15.02 11.05 8.52 - -
ApiQ 2 uniform 128 8.25 6.71 - - - 12.04 9.13 - - -
EfficientQAT 2 uniform 128 7.19 6.08 4.61 9.80 6.38 8.79 7.75 6.48 13.22 9.53
AQLM 2 vector 2x8 7.24 6.06 4.49 - - 8.96 7.80 6.36 - -
AQLM 2 vector 1x16 6.34 5.59 4.06 7.76 5.10 8.06 7.27 6.04 10.89 8.31
QuIP# 2 vector - 6.66 5.74 4.16 - - 8.35 7.45 6.12 - -
OmniQ 2 uniform 64 9.62 7.56 6.11 - - 12.72 10.05 7.68 - -
ApiQ 2 uniform 64 7.59 6.44 - - - 10.56 8.92 - - -
EfficientQAT 2 uniform 64 6.86 5.96 4.52 9.41 6.07 8.50 7.59 6.38 12.77 9.23

### v3

@@ Abstract
EfficientQAT: Efficient Quantization-Aware Training
for Large Language Models
Mengzhao Chen1,2, Wenqi Shao†2, Peng Xu1,2, Jiahao Wang1,2, Peng Gao2, Kaipeng Zhang2, Ping Luo†1 1The University of Hong Kong 2Shanghai AI Laboratory
footnotetext: Corresponding authors: shaowenqi@pjlab.org.cn; pluo@cs.hku.hk
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Touvron et al., 2023; Bubeck et al., 2023; Chiang et al., 2023; Xu et al., 2023a; Ying et al., 2024) have demonstrated impressive capabilities in diverse language tasks such as reasoning (Clark et al., 2018, 2019; Zellers et al., 2019), cognitive processing (Fu et al., 2023; Xu et al., 2023a), and agent-based applications (Qin et al., 2023a, b). However, these models are characterized by their extensive parameters, which pose significant challenges for memory footprint and bandwidth (Kim et al., 2023b; Xu et al., 2024a).
(a) 2-bit quantization comparisons
(b) Q-PEFT comparisons
Figure 1: (a) EfficientQAT significantly surpasses existing uniform quantization methods, and is either superior to or comparable with vector quantization techniques. (b) EfficientQAT markedly outperforms existing Q-PEFT methods.
@@
To address these challenges, we introduce a novel quantization-aware training framework called EfficientQAT. This framework combines the advantages of fully trainable parameters and end-to-end training, similar to native QAT (Ma et al., 2024), while maintaining the training efficiency of PTQ (Cheng et al., 2023; Shao et al., 2023) and Q-PEFT (Xu et al., 2023b). EfficientQAT introduces block-wise training of all parameters (Block-AP) to enhance the optimizable space and mitigate quantization accuracy loss. Block-AP sequentially trains all parameters, including original weights and quantization parameters (step sizes and zero points), within each transformer block. Several works have been developed based on block-wise reconstruction. However, previous approaches focus on designing additional trainable parameters, such as clipping thresholds for OmniQuant (Shao et al., 2023), weight rounding for AutoRound (Cheng et al., 2023) and BRECQ (Li et al., 2021), or LoRA (Hu et al., 2021) parameters for CBQ (Ding et al., 2023). Our Block-AP is the first to directly train all parameters during block-wise reconstruction, achieving superior performance compared to previous methods (see Table 5). Block-AP successfully demonstrates that complex trainable parameter design is unnecessary for effective block-wise reconstruction in LLMs quantization. Furthermore, we introduce end-to-end training of quantization parameters (E2E-QP) to account for inter-block interactions. E2E-QP keeps the quantized weights fixed and trains only the quantization parameters (step sizes) end-to-end.
Thanks to the integration of the proposed Block-AP and E2E-QP, EfficientQAT characterizes itself as a fast-converging, memory-efficient, and high-performing quantization technique. For instance, EfficientQAT can obtain a 2-bit Llama-2-70B model on a single A100-80GB GPU in just 41 hours, with less than 3 points accuracy degradation on 5 zero-shot common-sense tasks compared to its full-precision counterpart (69.48 vs. 72.41). We also evaluate EfficientQAT across scenarios involving model compression and instruction-tuning. In model compression, as illustrated in Figure 1(a), EfficientQAT significantly outperforms existing uniform quantization methods by approximately 5 points on accuracy in the challenging 2-bit quantization setting. In terms of instruction tuning, as shown in Figure 1(b), EfficientQAT consistently outperforms existing Q-PEFT methods, including QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and PEQA (Kim et al., 2023a). For instance, EfficientQAT surpasses PEQA (Kim et al., 2023a) with 4.5 points MMLU accuracy when fine-tuning with Alpaca dataset.
@@ 2 Related Works
Post-Training Quantization of LLMs. PTQ is a pivotal technique for accelerating and deploying LLMs. Quantization approaches generally fall into two categories: weight-only quantization (Frantar et al., 2022; Dettmers et al., 2023b; Lee et al., 2023a; Kim et al., 2023b) and weight-activation quantization (Xiao et al., 2023; Liu et al., 2023c; Wei et al., 2022, 2023; Yuan et al., 2023; Zhao et al., 2023; Ashkboos et al., 2023; Li et al., 2023a; Ashkboos et al., 2024). Weight-only quantization focuses on compressing weights into low-bit formats, reducing memory demands and enhancing the efficiency of memory-bounded computations in LLMs (Lin et al., 2024; Yuan et al., 2024). Conversely, weight-activation quantization compresses both weights and activations, thus further decreasing the overhead associated with matrix multiplications (Lin et al., 2024). Recent advancements in weight-only quantization include the introduction of vector quantization methods by QUIP#Tseng et al. (2024) and AQLMEgiazarian et al. (2024). These methods have shown promising performance but also introduce significant overhead (Gong et al., 2024). Our research continues to explore uniform quantization, which is preferred for its compatibility with hardware implementations.
@@ 3.1 Method Overview
In this section, we introduce EfficientQAT, a novel quantization-aware training framework for LLMs that enhances memory efficiency. As illustrated in Figure 2, traditional QAT approaches train the weights $\mathbf{W}$ and quantization parameters $s$ (step sizes) and $z$ (zero points) simultaneously in an end-to-end manner, which significantly increases the memory requirements due to the large number of parameters involved. To address this issue, EfficientQAT adopts a two-stage strategy: block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). In the Block-AP phase, model parameters and quantization parameters are trained block-by-block using reconstruction loss, which not only allows for precise calibration with full training but also reduces memory consumption (Li et al., 2021; Shao et al., 2023) by block-wise training. Following this, the E2E-QP phase fixes the quantized weights and trains the step sizes exclusively on target datasets, thus achieving inter-block interaction in a memory-efficient way. Details on Block-AP and E2E-QP are further described in Sections 3.2 and 3.3, respectively.
@@ 3.3 End-to-End Training of Quantization Parameters
Table 1: Llama 2 & 3 average zero-shot accuracy on 5 common-sense reasoning tasks ($\uparrow$). "-" indicates the result is unreachable in the public papers.
Method Bits Group 2-7 2-13 2-70 3-8 3-70
FP16 16 - 64.86 67.81 72.41 68.58 75.33
RTN 3 128 62.06 65.77 70.83 58.72 65.29
GPTQ 3 128 62.48 66.18 71.47 60.58 71.28
AWQ 3 128 62.82 66.14 71.41 64.82 73.65
OmniQ 3 128 62.42 66.18 71.07 64.09 71.90
AutoRound 3 128 63.72 66.68 71.24 - -
QuIP# 3 - 63.52 66.26 72.13 - -
EfficientQAT 3 128 64.02 67.28 71.76 67.35 73.96
OmniQ 2 128 46.98 53.56 54.87 52.66 60.06
AutoRound 2 128 54.50 60.72 67.70 - -
EfficientQAT 2 128 59.50 63.88 68.93 59.37 67.57
AQLM 2 2x8 57.61 62.22 69.85 - -
QuIP# 2 - 60.61 64.44 70.91 - -
DB-LLM 2 64 56.93 61.61 68.01 51.74 -
EfficientQAT 2 64 60.14 64.48 69.48 60.76 67.89
@@ 4.1 EfficientQAT for LLMs Quantization
PTQ Baseline. We compare our results with PTQ methods from uniform quantization such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), ApiQ (Liao and Monz, 2024) and AutoRound (Cheng et al., 2023), and vector quantization including QuIP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024). Note that if a result is the best of uniform quantization, we set it to bold.
Accuracy results. We evaluate the zero-shot accuracy on five common-sense reasoning tasks using the v0.4.2 lm-evaluation-harness***https://github.com/EleutherAI/lm-evaluation-harness. These tasks include WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), Arc-Easy (Clark et al., 2018), and Arc-Challenge (Clark et al., 2018). Table 1 shows that the proposed EfficientQAT significantly outperforms previous methods for uniform quantization across the Llama-2 and Llama-3 model families, as well as in both 2-bit and 3-bit quantization settings. The performance gains are particularly notable in extremely low-bit quantization, such as 2-bit. For instance, EfficientQAT achieves a +3.26% accuracy improvement over AWQ in w3g128 quantization with Llama-3-8B. Moreover, EfficientQAT surpasses DB-LLM by +9.02% accuracy in w2g64 quantization. In comparison to vector quantization, our results show that EfficientQAT outperforms QuIP#Tseng et al. (2024) in 3-bit quantization, but underperforms in 2-bit scenarios. However, direct comparisons between uniform quantization methods (such as EfficientQAT) and vector quantization methods (such as QuIP#) can be misleading due to fundamental differences in their approaches. Vector quantization often achieves better results at very low bit-widths through complex codebook designs, but this comes at the cost of reduced generalization and deployment flexibility. For instance, EfficientQAT supports both weight and activation quantization, while vector quantization methods are typically limited to weight-only quantization. Furthermore, a recent study, PrefixQuantChen et al. (2024b), demonstrates that EfficientQAT improves state-of-the-art weight-activation quantization methods by nearly 0.3 perplexity.
Perplexity results. We also evaluate perplexity on Wikitext2 and C4 using a 2048 context length, following prior studies (Frantar et al., 2022; Shao et al., 2023). The results align with the accuracy comparison, as EfficientQAT consistently achieves lower perplexity across the Llama-2 and Llama-3 model families in both 2-bit and 3-bit quantization. Notably, the benefits are more pronounced in Llama-3 models, which face greater challenges in quantization (Huang et al., 2024). For example, EfficientQAT reduces perplexity by 0.37 and 4.19 points compared to DB-LLM in Llama-2-7B and Llama-3-8B, respectively.
How model size and training tokens affect quantization error. Recent scaling laws for PTQ Kumar et al. (2024); Ouyang et al. (2024) show that quantization error increases with the number of training tokens and decreases as model size grows. Our results in Table 1 and Table 2 are consistent with these PTQ scaling laws. Additionally, the absolute benefit of our proposed method is more pronounced in smaller models, as they experience greater performance degradation from quantization. For example, DB-LLM loses 7.93 accuracy points with W2G64 on Llama-2-7B, but only 4.40 on Llama-2-70B. As a result, the improvement of EfficientQAT over DB-LLM decreases from 3.21 on Llama-2-7B to 1.47 on Llama-2-70B. However, when we use the relative gain metric $\frac{\text{EfficientQAT}-\text{DBLLM}}{\text{FP16}-\text{DBLLM}}$, EfficientQAT reduces quantization error by 40% for Llama-2-7B and 33% for Llama-2-70B. The relative gain metric demonstrates the effectiveness of proposed EfficientQAT across different model sizes.